What is the y-intercept of the function, f(x)=2x^2+4x-6?
Answer by venugopalramana(3286)

Y INTERCEPT IS DETERMINED BY THE POINTS WHERE THE CURVE INTERSECTS X AXIS.HENCE THIS IS
OBTAINED BY EQUATING THE EQN. TO ZERO AND FINDING VALUES OF X SATISFYING THE EQN.
What is the y-intercept of the function, y=f(x)=2x^2+4x-6?=0
x^2+2x-3=0
x^2+3x-x-3=0
x(x+3)-1(x+3)=0
(x+3)(x-1)=0
x=-3 and x=1 ARE THE 2 Y INTERCEPTS.